determine the mass of carbon dioxide produced during the complete combustion of 2.31g of this hydrocarbon fuel (ethanol)


1
Expert's answer
2021-01-11T03:53:44-0500

C2H5OH+3O2=3H2O+2CO2 

"n(C2H5OH)=\frac{m}{M}=\frac{2.31}{46,069}=0.05"

"\frac{0.05}{1}=\frac{x}{2}"

x=0.1

"m(CO2)=n\times M=0.1\times44.0095=4.4"